# Global parameterizations of $\pi\pi$ scattering with dispersive constraints: beyond the S0 wave

### v1

@@ III.4.2 The three G0-wave Global Fits
III.5 S2-wave parameterization
The S2 wave does not have much structure as no resonances appear there since it is a repulsive channel. Its CFD description in [54] was fairly good up to 1.42 GeV, but here we want to obtain a global fit up to the last available data point, almost at 2.1 GeV. In addition, we will allow its inelasticity to start at $\hat{s}=(0.915\,{\mathrm{\,Ge\kern-1.00006ptV}})^{2}$, which is just a convenient phenomenological value, with no other physical meaning than the approximate energy where the inelasticity seems to open in Solution B of [22]. This adds more flexibility compared to [54] where the inelasticity opened up at $1.05\,{\mathrm{\,Ge\kern-1.00006ptV}}$. Lowering the inelasticity effective threshold and extending the parameterization to higher energies require some minor changes and one more parameter compared to the CFD parameterization in [54], but our global parameterization remains rather simple and easy to implement.
@@ III.5 S2-wave parameterization
where, once again, the prime denotes the derivative with respect to $s$ and therefore,
@@
For the elasticity, we set $\eta_{0}^{(2)}(s)=1$ for $s$ below $\hat{s}=(0.915\,{\mathrm{\,Ge\kern-1.00006ptV}})^{2}$ and above, we use the empirical fit
@@
The elasticity data is so poor and scarce that one parameter is enough to describe them. Since no resonances are known below 2.1 GeV, above the last data point at 1.4 GeV, we just show an educated extrapolation of the elasticity up to 2.1 GeV, because that is the energy up to where data on the phase shift exists. Note, however, that we will only use this wave up to 1.62 GeV as input in our dispersive representation.
Figure 8: S2-wave phase shift (top) and elasticity (bottom). We show our unconstrained and dispersively constrained Global Fit I together with the CFD from [54]. As explained in the text, the data used in the fit above 0.9 GeV are those with solid symbols from Cohen et al. [23], OPEDP from Durusoy et al. [21] and the A solutions from Losty et al. [22] and Hoogland et al. [27]. The other data are shown for completeness. In the top panel, the black vertical line at 0.85 GeV marks the matching point with the high-energy parameterization. In the bottom one, the red vertical line stands at 0.915 GeV, where we allow the inelasticity to open up instead of 1.05 GeV (blue vertical line) used in the CFD.
@@ III.5.1 S2-wave Global Fit and parameters
In Fig. 8, we show the phase shift and elasticity of our constrained Global Fit I, together with all available scattering data. However, the input we use in our fit is the CFD below 0.9 GeV and only the data represented by solid symbols above 0.9 GeV and the $a_{0}^{(2)}$ and $b_{0}^{(2)}$ threshold parameters in Table 10 in Appendix A.
Regarding the phase shifts, Cohen et al. [23] only provide one set of data, whereas Durusoy et al. [21], Losty et al. [22], and Hoogland et al. [27] each provide two sets of data, which correspond to slightly different methods of analyses that overlap within uncertainties. We show all sets for completeness but fit only one per experiment (the solid symbols in Fig. 8). For [21] and [27] we have chosen to fit only the set that includes a Dürr-Pilkuhn [91] form factor, for consistency with the data from [23]. This is the most conservative choice since it has somewhat larger uncertainties. Concerning [22], we fit Solution A, discarding Solution B for two reasons. First, it has larger uncertainties than the rest of the data and including them or not in the S2 fit is almost irrelevant. The fit is still compatible with Solution B as long as we allow the inelasticity to open up at 0.915 GeV. Second, we will see that for the G2 wave, and despite these huge uncertainties, Solution B is incompatible with the rest of the data. We will keep these choices of data input for all the $I=2$ waves.
The Global Fit I phase shift is compatible with the CFD parameterization in [54], valid only up to 1.42 GeV. However, the new phase shift, which fits data up to 2.1 GeV, reaches a minimum near 1.3 GeV and grows towards zero at higher energies.
A significant difference with the CFD is seen in the inelasticity, which we now allow to start from 0.915 GeV, instead of 1.05 GeV, as for the CFD. We reiterate that this value has no other physical meaning than the energy where the data starts showing some non-vanishing inelasticity. We prefer our new global parameterization in Eq. (45) since it is continuous, differentiable, and more consistent with the two data points from Cohen et al. [23]. As explained above, the Solution B data from Losty et al. [22], even if not included in the fit, are quite compatible with it due to their large uncertainties. Beyond 1.5 GeV, our elasticity is just a naive extrapolation, since no data exist beyond that point. Durusoy et al. [21] used a linear extrapolation in $\sqrt{s}$ of the two data points of Cohen et al. [23]. Nevertheless, they say that their phase shifts are not very sensitive to different elasticity guesses. The S2-wave Global Fit I parameters after using the dispersion relations as constraints are given in Table 5.
@@ III.5.2 The three S2-wave Global Fits
In contrast to the S0, P, D0, and F waves, each $I=2$ wave has only one data set to fit from the start. Therefore we expect the three constrained Global Fits to the $I=2$ waves to differ very little among themselves since their separation is an effect induced indirectly from the other waves used as input in the dispersion relations.
@@
III.6 D2-wave parameterization
Except close to the $\pi\pi$ threshold, this is a repulsive wave, with no resonances and little structure, even smaller than the S2. The CFD parameterization in [54] only reached $\sim 1.4\,$GeV. Here, we extend the parameterization up to 2.1 GeV, where a few data on its phase shift exist. It will have two pieces with continuous and differentiable matching at 1.4 GeV.
The phase shift parameterization will be very similar to that of the S2, except for the appropriate angular momentum factors and another factor to ensure a positive scattering length (confirmed by sum rules and Chiral Perturbation Theory calculations), despite the wave being mostly repulsive.
@@ III.6 D2-wave parameterization
With these definitions, both the parameterization and its derivative are continuous at $s_{m}$.
@@
III.6.1 D2 Global Fit and parameters
For this wave, we include in the fit the values of the threshold parameters $a^{(2)}_{2}$ and $b^{(2)}_{2}$ obtained from the CFD parameterization in [54], given in Table 10 here. In addition, we fit the CFD results below 0.915 GeV, as well as the data on the phase shift that we show in Fig. 10. It should be noted that this was the wave where the dispersive constraints demanded the largest change between the Unconstrained and Constrained Fit to Data (UFD versus CFD in [54]). This is a strong reason to use the CFD as input and we see that, up to energies around 0.95 GeV, the curves lie above most of the data.
Figure 10: D2-wave phase shift. We show our unconstrained and dispersively constrained Global Fit I, together with the dispersively constrained CFD parameterization from [54]. As explained in the text, the data used in the fit above 0.9 GeV are those with solid symbols. The other data are shown for completeness. The black vertical line at 0.85 GeV marks the matching point with the high-energy part. All experiments either do not observe or assume that there is no inelasticity. Data references as in Fig. 8.
Compared to the CFD in [54], we now extend our parameterization to 2.1 GeV, since there is still a data point at this energy. When two consistent data sets from the same experiment are available, as in Durusoy et al. [21] and Hoogland et al. [27], we only fit one set, chosen with the same criteria explained for the S2 wave. Once again the data of Losty et al. [22] lie systematically below and tend to have much larger uncertainties than other experiments in the 0.8 to 1.2 GeV region.
The D2-wave parameters are given in Table 6.
III.6.2 The three D2-wave Global Fits
III.7 G2-wave parameterization
Once again, this is a repulsive wave, very small in the whole energy range of interest. In the dispersive analysis of [54], it was neglected below 1.42 GeV, based on previous estimations [73] that studied its negligible impact on dispersive calculations within this range. However, there are a few data points available up to almost 2.1 GeV, and hence we now provide a global parameterization. In addition, its role is not so small in the two FDRs we study above 1.4 GeV here.
In particular, the parameterization in [73] was little more than an order of magnitude estimate above 1 GeV, assuming its phase shift to be zero below. No attempt was made to describe the threshold parameters.
@@ III.7 G2-wave parameterization
We will use a parameterization of the phase shift rather similar to that of the D2 wave. Namely, for $s^{1/2}\leq s_{m}^{1/2}=0.85\,{\mathrm{\,Ge\kern-1.00006ptV}}$ we use a conformal expansion:
@@
and both the parameterization and its derivative with respect to $s$ are continuous at the matching point.
@@
III.7.1 G2 Global Fit and parameters
Three different experimental collaborations, Cohen et al. [23], Durusoy et al. [21], and Losty et al. [22] provide measurements of the phase shift for this wave, shown in Fig. 12. The set from Cohen et al. [23] only has two data points at 1.2 and 1.4 GeV, but they are remarkably compatible with the data from Durusoy et al. [21], which has nine points spanning from 0.5 to 2.1 GeV, and hence, dominates our Global Fit. Following the same criteria we have already used for other $I=2$ waves, we have fit the data from  [23] and only one of the two analyses from Durusoy et al. (OPEDP) [21]. Moreover, although only Solution B of Losty et al. [22] considers the G-wave, it has not been included in the fit, because it is incompatible with the other two sets above 0.9 GeV, despite its much larger uncertainties.
@@ IV Dispersion relations
Since a two-body scattering amplitude $F(s,t)$ depends on two independent variables, in order to apply Cauchy’s integral theorem it is convenient to either fix $t$ or integrate it to define partial waves. The first option gives rise to fixed-$t$ dispersion relations and the second to partial-wave dispersion relations. Next, we will discuss both the fulfillment and the use as constraints of the two most common instances of both types of dispersion relations: forward dispersion relations and Roy-like equations.
@@ IV.1.1 Definitions
Let us first discuss the fixed $t=0$ dispersion relations, also known as forward dispersion relations. They are very relevant for three reasons. First, their applicability can be extended, in principle, to any value of $s$. Second, the optical theorem relates the imaginary part of the forward amplitude to the total cross-section, for which data are easier to obtain. Third, when the combination of amplitudes is chosen conveniently, the integrands are all positive and the resulting uncertainties are generically small.
@@
This one is not definite positive and its uncertainties will be generically larger.
Note that the integral of dispersion relations extends up to infinity. At high energies, above a certain matching energy, we use the Regge parameterizations of $\pi\pi$ amplitudes obtained from $\pi\pi$, $\pi N$ and $NN$ cross sections in [95] and upgraded and updated in [54] and the review [45]. In a later subsection, we will discuss the matching with these Regge parameterizations.
In what follows we will first check how well the FDRs are satisfied with the partial-wave fits to the data discussed above, and then, we will use them as constraints on the fits.
Figure 14: Fulfillment of forward dispersion relations before (left column) and after (right column) they have been imposed as constraints on the Global Fit I. The continuous line corresponds to the evaluation of $\textrm{Re}F^{i}$ with the dispersive integral and using the global parameterizations as input, whereas the dashed line is calculated directly from the global parameterizations. The red band is the error in their difference, that we attach to the “direct” one for illustration. The vertical blue line at 1 GeV separates the two energy regions, for which we provide separated averaged quadratic distances $\bar{d}^{2}$. The unconstrained Global Fit I only satisfies well the FDRs below 1 GeV ($\bar{d}^{2}<1$), but is rather inconsistent above ($\bar{d}^{2}>>1$). In contrast, the constrained Global Fit I satisfies very well the FDRs in the two energy regions.
@@ IV.1.2 FDRs as checks
In [70, 73, 71, 72, 54] the three FDRs in Eqs. (IV.1.1) and (57) were imposed as constraints of phenomenological fits to data, to be satisfied within uncertainties to obtain the CFD parameterizations. Moreover, in [54] Roy and GKPY equations—to be explained below—were imposed on the CFD too. Since our global parameterizations mimic the CFD sets up to 0.9 GeV, it is no surprise that the FDRs are well satisfied by our Global Fits in that region, even without using them as constraints. This can be seen in the left panels of Fig. 14, where we show the fulfillment of FDRs before we impose them as constraints in our Global Fits to data. Note that in Fig. 14 we show as dashed lines the parts of the FDRs that are calculated directly from the parameterizations and as continuous lines the same quantities obtained from the integral representation, which we call “dispersive”. In theory (in the isospin limit, without $4\pi$ inelasticity, etc…), these two lines should agree, and their difference should be zero. However, since ours is a data analysis, we only expect them to agree within uncertainties. The uncertainties of the “direct” and “dispersive” parts are correlated, as they are both calculated from the same parameterizations. Such correlations cancel out to a large degree in their difference. Thus, we calculate the uncertainty band of their difference and, for illustrative purposes, we attach it to the direct part. Thus, we aim at a fit such that the continuous line would fall within that red uncertainty band. From the left panels of Fig. 14 it is clear that the three FDRs are not well satisfied above the 1 GeV region before we impose them as constraints into our Global Fits.
@@
For the above reasons, we impose the FDRs as constraints of our Global Fits. Note that we first impose the FDRs, because our global parameterizations here mimic the CFD below 0.9 GeV and only deviate significantly from them above that energy. The applicability range of Roy and GKPY equations extends up to approximately 1.11.11.1\,1.1GeV and were already imposed in the CFD in [70, 73, 71, 72]. Since they have relatively large uncertainties above 0.9 GeV compared to FDRs, they are still fairly well satisfied with only a minor deviation around 1.1 GeV for the S2 GKPY equation. Thus, we first impose the FDRs and, on a second step discussed in a subsection below, we also impose the Roy and GKPY equations, which only produce a minor modification to reach our final Global Fits. As in previous sections, we only illustrate this procedure in detail for Global Fit I, and we just discuss the results for Global Fit II and III later on.
@@
It is important to emphasize that the error bands in the 0.9 to 1.4 GeV region of the FDRs for the Global Fits obtained in this work are smaller than those of previous works [54, 68]. This is clearly seen in Fig. 15, where we plot the $\Delta^{i}(s),i=00,0+,I_{t}=1$ differences defined in Eqs. (IV.1.1) and (57) both for our Global Fit I and the CFD parameterizations in [54]. The new bands have become visibly smaller, mostly due to our much-improved P wave above the $\pi\omega$ threshold, and to a lesser extent, to the slight improvement in the D0 wave. The largest uncertainty reduction occurs in $\Delta^{0+}$, which is dominated by the P wave, followed by the $\Delta^{I_{t}=1}$, where the P wave also has a large contribution. The $\Delta^{00}$ uncertainty band, where the P wave does not contribute, is only slightly smaller than in previous works. For this reason, the FDRs in the 0.9 to 1.4 GeV region, now extended to 1.6 GeV for two of them, have become even more stringent constraints than in previous works, and their compliance is a more notable feature.
@@
The situation regarding the $F^{I_{t}=1}$ FDR is more telling, since Global Fits II, and III satisfy the $I_{t}=1$ FDR in a clearly less uniform way than Global Fit I. This is shown in Fig. 16, where we plot the $\Delta^{I_{t}=1}$ differences and their uncertainties for the three Global Fits. Although the average $\bar{d}_{I_{t}=1}^{2}<1$, both below and above 1 GeV, only the Global Fit 1 has $\bar{d}_{I_{t}=1}^{2}<1$ everywhere. In contrast, for Global Fits II and III, $\Delta^{I_{t}=1}$ lies beyond one deviation from zero in two regions wider than 100 MeV, around 1 GeV and 1.5 GeV. As it can be seen in Table 8, for these two constrained Global Fits, the $F^{I_{t}=1}$ FDR yields $\bar{d}^{2}>1.5$ between $0.93$ and $1.06{\mathrm{\,Ge\kern-1.00006ptV}}$. In addition, Global Fit III yields $\bar{d}^{2}\simeq 1.5$ from $1.46$ to $1.56{\mathrm{\,Ge\kern-1.00006ptV}}$. The $0.93-1.06{\mathrm{\,Ge\kern-1.00006ptV}}$ region is of particular interest because it is where most waves become inelastic. Actually, these small discrepancies in Global Fits II and III can be amended if we set the opening of their D0-wave inelasticity below the $K\bar{K}$ threshold, as done for Global Fit I. However, as explained in previous sections, this is not how the data for Solutions II and III were obtained. Hence, in this respect, Global Fits II and III are, once again, slightly disfavored compared to Global Fit I, and in particular, the opening of the D0-wave inelasticity below 1 GeV is strongly favored.

### v2

@@ III.4.2 The three G0-wave Global Fits
III.5 S2-wave parametrization
The S2 wave does not have much structure as no resonances appear there since it is a repulsive channel. Its CFD description in [54] was fairly good up to 1.42 GeV, but here we want to obtain a global fit up to the last available data point, almost at 2.1 GeV. In addition, we will allow its inelasticity to start at $\hat{s}=(0.915\,{\mathrm{\,Ge\kern-1.00006ptV}})^{2}$, which is just a convenient phenomenological value, with no other physical meaning than the approximate energy where the inelasticity seems to open in Solution B of [22]. This adds more flexibility compared to [54], where the inelasticity opened up at $1.05\,{\mathrm{\,Ge\kern-1.00006ptV}}$. Lowering the inelasticity effective threshold and extending the parametrization to higher energies require some minor changes and one more parameter compared to the CFD parametrization in [54], but our global parametrization remains rather simple and easy to implement.
@@ III.5 S2-wave parametrization
where, once again, the prime denotes the derivative with respect to $s$ and, therefore,
@@
For the elasticity, we set $\eta_{0}^{(2)}(s)=1$ for $s$ below $\hat{s}=(0.915\,{\mathrm{\,Ge\kern-1.00006ptV}})^{2}$ and above we use the empirical fit
@@
The elasticity data is so poor and scarce that one parameter is enough to describe them. Since no resonances are known below 2.1 GeV, above the last data point at 1.4 GeV, we will just show an educated extrapolation of the elasticity up to 2.1 GeV because that is the energy up to where data on the phase shift exists. Note, however, that we will only use this wave up to 1.62 GeV as input in our dispersive representation.
Figure 8: S2-wave phase shift (top) and elasticity (bottom). We show our unconstrained and dispersively constrained Global Fit I together with the CFD from [54]. As explained in the text, the data used in the fit above 0.9 GeV are those with solid symbols from Cohen et al. [23], OPEDP from Durusoy et al. [21] and the A solutions from Losty et al. [22] and Hoogland et al. [27]. The other data are shown for completeness. In the top panel, the black vertical line at 0.85 GeV marks the matching point with the high-energy parametrization. In the bottom one, the red vertical line stands at 0.915 GeV, where we allow the inelasticity to open up instead of 1.05 GeV (blue vertical line) used in the CFD.
@@ III.5.1 S2-wave Global Fit and parameters
In Fig. 8, we show the phase shift and elasticity of our constrained Global Fit I, together with all available scattering data. However, the input we use in our fit is the CFD below 0.9 GeV, the data represented by solid symbols above 0.9 GeV and the $a_{0}^{(2)}$ and $b_{0}^{(2)}$ threshold parameters in Table 10 in Appendix A.
Regarding the phase shifts, Cohen et al. [23] only provide one set of data, whereas Durusoy et al. [21], Losty et al. [22], and Hoogland et al. [27] each provide two sets of data, which correspond to slightly different methods of analyses that overlap within uncertainties. We show all sets for completeness but fit only one per experiment (the solid symbols in Fig. 8). For [21] and [27] we have chosen to fit only the set that includes a Dürr-Pilkuhn [95] form factor for consistency with the data from [23]. This is the most conservative choice since it has somewhat larger uncertainties. Concerning [22], we fit Solution A, discarding Solution B for two reasons. First, it has larger uncertainties than the rest of the data, and including them or not in the S2 fit is almost irrelevant. The fit is still compatible with Solution B as long as we allow the inelasticity to open up at 0.915 GeV. Second, we will see that for the G2 wave, and despite these huge uncertainties, Solution B is incompatible with the rest of the data. We will keep these choices of data input for all the $I=2$ waves.
The Global Fit I phase shift is compatible with the CFD parametrization in [54], valid only up to 1.42 GeV. However, the new phase shift, which fits data up to 2.1 GeV, reaches a minimum near 1.3 GeV and grows toward zero at higher energies.
A significant difference with the CFD is seen in the inelasticity, which we now allow to start from 0.915 GeV instead of 1.05 GeV, as for the CFD. We reiterate that this value has no other physical meaning than the energy where the data starts showing some non-vanishing inelasticity. We prefer our new global parametrization in Eq. (45) since it is continuous, differentiable, and more consistent with the two data points from Cohen et al. [23]. As explained above, the Solution B data from Losty et al. [22], even if not included in the fit, are quite compatible with it due to their large uncertainties. Beyond 1.5 GeV, our elasticity is just a naive extrapolation since no data exist beyond that point. Durusoy et al. [21] used a linear extrapolation in $\sqrt{s}$ of the two data points of Cohen et al. [23]. Nevertheless, they say that their phase shifts are not very sensitive to different elasticity guesses. The S2-wave Global Fit I parameters, after using the dispersion relations as constraints are given in Table 5.
@@ III.5.2 The three S2-wave Global Fits
In contrast to the S0, P, D0, and F waves, each $I=2$ wave has only one data set to fit from the start. Therefore, we expect the three constrained Global Fits to the $I=2$ waves to differ very little among themselves since their separation is an effect induced indirectly from the other waves used as input in the dispersion relations.
@@
III.6 D2-wave parametrization
Except close to the $\pi\pi$ threshold, this is a repulsive wave, with no resonances and little structure, even smaller than the S2. The CFD parametrization in [54] only reached $\sim 1.4\,$GeV. Here, we extend the parametrization up to 2.1 GeV, where a few data on its phase shift exist. It will have two pieces with continuous and differentiable matching at 1.4 GeV.
The phase shift parametrization will be very similar to that of the S2, except for the appropriate angular momentum factors and another factor to ensure a positive scattering length (confirmed by sum rules and Chiral Perturbation Theory calculations), despite the wave being mostly repulsive.
@@ III.6 D2-wave parametrization
With these definitions, both the parametrization and its derivative are continuous at $s_{m}$.
Figure 10: D2-wave phase shift. We show our unconstrained and dispersively constrained Global Fit I, together with the dispersively constrained CFD parametrization from [54]. As explained in the text, the data used in the fit above 0.9 GeV are those with solid symbols. The other data are shown for completeness. The black vertical line at 0.85 GeV marks the matching point with the high-energy part. All experiments either do not observe or assume that there is no inelasticity. Data references as in Fig. 8.
@@
III.6.1 D2-wave Global Fit and parameters
For this wave, we include in the fit the values of the threshold parameters $a^{(2)}_{2}$ and $b^{(2)}_{2}$ obtained from the CFD parametrization in [54], given in Table 10 here. In addition, we fit the CFD results below 0.915 GeV, as well as the data on the phase shift that we show in Fig. 10. It should be noted that this was the wave where the dispersive constraints demanded the largest change between the Unconstrained and Constrained Fit to Data (UFD versus CFD in [54]). This is a strong reason to use the CFD as input, and we see that, up to energies around 0.95 GeV, the curves lie above most of the data.
@@ III.6.1 D2-wave Global Fit and parameters
Compared to the CFD in [54], we now extend our parametrization to 2.1 GeV, since there is still a data point at this energy. When two consistent data sets from the same experiment are available, as in Durusoy et al. [21] and Hoogland et al. [27], we only fit one set, chosen with the same criteria explained for the S2 wave. Once again, the data of Losty et al. [22] lie systematically below and tend to have much larger uncertainties than other experiments in the 0.8 to 1.2 GeV region.
The D2-wave parameters are given in Table 6.
III.6.2 The three D2-wave Global Fits
@@
III.7 G2-wave parametrization
Once again, this is a repulsive wave, very small in the whole energy range of interest. In the dispersive analysis of [54], it was neglected below 1.42 GeV, based on previous estimations [73] that studied its negligible impact on dispersive calculations within this range. However, there are a few data points available up to almost 2.1 GeV, and hence, we now provide a global parametrization. In addition, its role is not so small in the two FDRs we study above 1.4 GeV here.
In particular, the parametrization in [73] was little more than an order of magnitude estimate above 1 GeV, assuming its phase shift to be zero below. No attempt was made to describe the threshold parameters.
We will use a parametrization of the phase shift rather similar to that of the D2 wave. Namely, for $s^{1/2}\leq s_{m}^{1/2}=0.85\,{\mathrm{\,Ge\kern-1.00006ptV}}$ we use a conformal expansion:
@@ III.7 G2-wave parametrization
and both the parametrization and its derivative with respect to $s$ are continuous at the matching point.
@@
III.7.1 G2-wave Global Fit and parameters
Three different experimental collaborations, Cohen et al. [23], Durusoy et al. [21], and Losty et al. [22] provide measurements of the phase shift for this wave, shown in Fig. 12. The set from Cohen et al. [23] only has two data points at 1.2 and 1.4 GeV, but they are remarkably compatible with the data from Durusoy et al. [21], which has nine points spanning from 0.5 to 2.1 GeV, and hence, dominates our Global Fit. Following the same criteria we have already used for other $I=2$ waves, we have fit the data from  [23] and only one of the two analyses from Durusoy et al. (OPEDP) [21]. Moreover, although only Solution B of Losty et al. [22] considers the G-wave, it has not been included in the fit because it is incompatible with the other two sets above 0.9 GeV, despite its much larger uncertainties.
@@ IV Dispersion relations
Since a two-body scattering amplitude $F(s,t)$ depends on two independent variables, in order to apply Cauchy’s integral theorem, it is convenient to either fix $t$ or integrate it to define partial waves. The first option gives rise to fixed-$t$ dispersion relations and the second to partial-wave dispersion relations. Next, we will discuss both the fulfillment and the use as constraints of the two most common instances of both types of dispersion relations: forward dispersion relations and Roy-like equations.
@@ IV.1.1 Definitions
Let us first discuss the fixed $t=0$ dispersion relations, also known as forward dispersion relations. They are very relevant for three reasons. First, their applicability can be extended, in principle, to any value of $s$. Second, the optical theorem relates the imaginary part of the forward amplitude to the total cross section, for which data are easier to obtain. Third, when the combination of amplitudes is chosen conveniently, the numerators of the integrands are all positive, and the resulting uncertainties are generically small.
@@
This one is not definite positive, and its uncertainties will be generically larger.
Note that the integral of dispersion relations extends up to infinity. At high energies, above a certain matching energy, we use the Regge parametrizations of $\pi\pi$ amplitudes obtained from $\pi\pi$, $\pi N$, and $NN$ cross sections in [99] and upgraded and updated in [54] and the review [45]. In a later subsection, we will discuss the matching with these Regge parametrizations.
In what follows, we will first check how well the FDRs are satisfied with the partial-wave fits to the data discussed above, and then we will use them as constraints on the fits.
Figure 14: Fulfillment of forward dispersion relations before (left column) and after (right column), they have been imposed as constraints on the Global Fit I. The continuous line corresponds to the evaluation of $\textrm{Re}F^{i}$ with the dispersive integral and using the global parametrizations as input, whereas the dashed line is calculated directly from the global parametrizations. The red band is the error in their difference, that we attach to the “direct” one for illustration. The vertical blue line at 1 GeV separates the two energy regions, for which we provide separated averaged quadratic distances $\bar{d}^{2}$. The unconstrained Global Fit I only satisfies well the FDRs below 1 GeV ($\bar{d}^{2}<1$), but is rather inconsistent above ($\bar{d}^{2}>>1$). In contrast, the constrained Global Fit I satisfies very well the FDRs in the two energy regions.
@@ IV.1.2 FDRs as checks
In [70, 73, 71, 72, 54] the three FDRs in Eqs. (IV.1.1) and (57) were imposed as constraints of phenomenological fits to data, to be satisfied within uncertainties to obtain the CFD parametrizations. Moreover, in [54] Roy and GKPY equations—to be explained below—were imposed on the CFD too. Since our global parametrizations mimic the CFD sets up to 0.9 GeV, it is no surprise that the FDRs are well satisfied by our Global Fits in that region, even without using them as constraints. This can be seen in the left panels of Fig. 14, where we show the fulfillment of FDRs before we impose them as constraints in our Global Fits to data. Note that in Fig. 14 we show as dashed lines the parts of the FDRs that are calculated directly from the parametrizations and as continuous lines the same quantities obtained from the integral representation, which we call “dispersive”. In theory (in the isospin limit, without $4\pi$ inelasticity, etc), these two lines should agree, and their difference should be zero. However, since ours is a data analysis, we only expect them to agree within uncertainties. The uncertainties of the “direct” and “dispersive” parts are correlated, as they are both calculated from the same parametrizations. Such correlations cancel out to a large degree in their difference. Thus, we calculate the uncertainty band of their difference and, for illustrative purposes, we attach it to the direct part. Thus, we aim at a fit such that the continuous line would fall within that red uncertainty band. From the left panels of Fig. 14 it is clear that the three FDRs are not well satisfied above the 1 GeV region before we impose them as constraints into our Global Fits.
@@
For the above reasons, we impose the FDRs as constraints of our Global Fits. Note that we first impose the FDRs, because our global parametrizations here mimic the CFD below 0.9 GeV and only deviate significantly from them above that energy. The applicability range of Roy and GKPY equations extends up to approximately 1.11.11.1\,1.1GeV and were already imposed in the CFD in [70, 73, 71, 72]. Since they have relatively large uncertainties above 0.9 GeV compared to FDRs, they are still fairly well satisfied with only a minor deviation around 1.1 GeV for the S2 GKPY equation. Thus, we first impose the FDRs and, on a second step discussed in a subsection below, we also impose the Roy and GKPY equations, which only produce a minor modification to reach our final Global Fits. As in previous sections, we only illustrate this procedure in detail for Global Fit I, and we just discuss the results for Global Fit II and III later on.
@@
It is important to emphasize that the error bands in the 0.9 to 1.4 GeV region of the FDRs for the Global Fits obtained in this work are smaller than those of previous works [54, 68]. This is clearly seen in Fig. 15, where we plot the $\Delta^{i}(s),i=00,0+,I_{t}=1$ differences defined in Eqs. (IV.1.1) and (57) both for our Global Fit I and the CFD parametrizations in [54]. The new bands have become visibly smaller, mostly due to our much-improved P wave above the $\pi\omega$ threshold, and to a lesser extent, to the slight improvement in the D0 wave. The largest uncertainty reduction occurs in $\Delta^{0+}$, which is dominated by the P wave, followed by the $\Delta^{I_{t}=1}$, where the P wave also has a large contribution. The $\Delta^{00}$ uncertainty band, where the P wave does not contribute, is only slightly smaller than in previous works. For this reason, the FDRs in the 0.9 to 1.4 GeV region, now extended to 1.6 GeV for two of them, have become even more stringent constraints than in previous works, and their compliance is a more notable feature.
@@
The situation regarding the $F^{I_{t}=1}$ FDR is more telling, since Global Fits II, and III satisfy the $I_{t}=1$ FDR in a clearly less uniform way than Global Fit I. This is shown in Fig. 16, where we plot the $\Delta^{I_{t}=1}$ differences and their uncertainties for the three Global Fits. Although the average $\bar{d}_{I_{t}=1}^{2}<1$, both below and above 1 GeV, only the Global Fit I has $\bar{d}_{I_{t}=1}^{2}<1$ everywhere. In contrast, for Global Fits II and III, $\Delta^{I_{t}=1}$ lies beyond one deviation from zero in two regions wider than 100 MeV, around 1 GeV and 1.5 GeV. As it can be seen in Table 8, for these two constrained Global Fits, the $F^{I_{t}=1}$ FDR yields $\bar{d}^{2}>1.5$ between $0.93$ and $1.06{\mathrm{\,Ge\kern-1.00006ptV}}$. In addition, Global Fit III yields $\bar{d}^{2}\simeq 1.5$ from $1.46$ to $1.56{\mathrm{\,Ge\kern-1.00006ptV}}$. The $0.93-1.06{\mathrm{\,Ge\kern-1.00006ptV}}$ region is of particular interest because it is where most waves become inelastic. Actually, these small discrepancies in Global Fits II and III can be amended if we set the opening of their D0-wave inelasticity below the $K\bar{K}$ threshold, as done for Global Fit I. However, as explained in previous sections, this is not how the data for Solutions II and III were obtained. Hence, in this respect, Global Fits II and III are, once again, slightly disfavored compared to Global Fit I, and in particular, the opening of the D0-wave inelasticity below 1 GeV is strongly favored.